# Multi-Agent Probabilistic Ensembles with Trajectory Sampling for Connected Autonomous Vehicles

### v1

@@ VI Experimental Settings and Numerical Results
In this section, we evaluate the performance of MA-PETS and demonstrate the superiority of our proposed algorithm over other several state-of-the-art MFRL methods, including PPO [8], DDPG [10], DQN [48], SAC [49]. Specifically, we run our experiments using the CAV simulation platform SMARTS [26] and select the “Unprotected Intersection” scenario in the closed “Figure-Eight” loop, which is a typical mixed autonomy traffic scenario as illustrated in Fig. 3, for the evaluation. We randomly deploy $I=8$ CAVs and a random number of HVs, where the former is controlled by our MA-PETS algorithm and the latter is controlled by the background environment in SMARTS. All CAVs will start from a one-way lane with an intersection, and drive circularly by passing through the intersection and avoiding collisions and congestion. Following the MDP defined in Section III, the corresponding MDP for CAVs in “Unprotected Intersection” is defined as below.
@@
  • Action: As each CAV is only controlled via the target velocity $v\in\mathbb{R}$ decided by itself, the action of vehicle $i$ can be represented as $a_{t}^{(i)}=v_{t}^{(i)}$.
  • Reward: The goal of CAVs is to maintain a maximum velocity based on no collision. Accordingly, the reward function $\mathcal{R}^{(i)}$ can be defined as
@@
(a) Aerial View
(b) Regional Enlarged View
Figure 3: The “Unprotected Intersection” scenario for Simulations.
TABLE II: List of Key Parameters for the Simulation.
Parameters Description Settings
Number of RL agents $I=8$
Communication range $d\in[0,200]$
Speed per vehicle at time-step $t$ $v_{t}\in\mathbb{R}$
@@
Target velocity per vehicle at $t$ $\bar{v}_{t}\in\mathbb{R}$
Speed of vehicles ahead and behind $v_{t,a},v_{t,e}\in\mathbb{R}$
Distance of vehicles ahead and behind $l_{t,a},l_{t,e}\in\mathbb{R}$
@@
Maximal length per episode $H=200$
No. of Ensembles for dynamics Model $B=5$
@@
In our setting, the number of episodes $K=15$, while the length of each episode is up to $H=200$ times-steps in case of no collision. We run 5 independent simulations. Besides, typical parameters are summarized in Table II.
@@
Firstly, we evaluate the performance of MA-PETS for CAVs with communication range $d=100$ and present the performance comparison with MFRL algorithms in Fig. 4. It can be observed from Fig. 4, that our algorithm MA-PETS significantly yields superior outcomes than the others in terms of both agility and safety. More importantly, MA-PETS converges at an apparently faster pace.
(a) Agility
(b) Safety
Figure 4: Comparison of utility in “Unprotected Intersection” scenario.
@@
Furthermore, we show the performance of MA-PETS with respect to different values of communication range $d$ by varying from 00 to $200$ in Fig. 5 and Table III. It can be observed from Table III that consistent with our intuition, the increase in communication range leads to a significant boost of the communications overheads. Meanwhile, as depicted in Fig. 5, it also benefits the learning efficiency of the CAVs is enhanced, thus greatly upgrading the agility and safety. On the other hand, Fig. 5 also unveils that when the communication range increases to a certain extent, a further increase of $d$ contributes trivially to the agility and safety of learned policies. In contrast, Table III indicates an exponential increase in the average communications overhead. In other words, it implies a certain trade-off between the learning performance and communication overheads. Consistent with the group regret bound detailed in Section V, we further investigate the numerical interplay between the minimum number of clique covers $\bar{\chi}\left(\mathcal{G}_{d,k}\right)$ and communication distance $d$ in $15$ independent trials. It can be observed from Fig. 6 that while the minimum number of clique covers remains constant beyond a specific communication range threshold, a generally inverse correlation exists between the minimum clique cover number and the communication radius, which aligns with both our theoretical proofs and the experimental findings presented in Fig. 5.
(a) Agility
(b) Safety
Figure 5: Performance comparison under different communication range $d$.
@@
Communication Distance $d$
00 $50$ $100$ $150$ $200$
Communication Overheads
00 $330.13$ $1,866$ $6,139.7$ $10,961.06$
@@ 00 $50$ $100$ $150$ $200$
To clarify the effect of the MPC horizon $w$ in MA-PETS, we perform supplementary experiments in “Figure Eight”. As depicted in Fig. 7, from a security point of view, extending the MPC horizon $w$ gradually enhances security. However, beyond a horizon length of $w=25$, further increases yield only marginal security improvements. In terms of agility, the choice of the MPC horizon $w$ is crucial for the performance of the system. In particular, a “too short” horizon introduces a substantial bias in the performance of MA-PETS, as it hampers the ability to make accurate long-term predictions due to the scarcity of time steps. Conversely, a “too long” horizon also results in a noticeable bias. This is due to the increased divergence of particles over extended periods, which reduces the correlation between the currently chosen action and the long-term expected reward. This bias escalates as the horizon $w$ increases. Hence choosing an excessively long horizon adversely affects performance.
(a) Agility
(b) Safety
Figure 7: Performance comparison under different horizon $w$ of MPC.
Our research also investigates the influence of the number of ensembles $B$ and the number of particles $P$ in MA-PETS. In terms of agility, Fig. 8 and Fig. 9 show the corresponding results respectively, which are derived from $15$ training episodes through $10$ independent simulation runs. It can be observed from Fig. 8 that along with the increase in $B$, the learning becomes more regularized and the performance improves. However, the performance improvement is no longer apparent for a sufficiently large $B$, and this improvement is more pronounced in more challenging and complex environments that require the learning of intricate dynamical models, leaving more scope for effective exploitation of the strategy without the use of model integration. In Fig. 9, superior performance can be reaped for a larger number of particles, due to the fact that more particles allow for a more accurate estimation of the reward for state-action trajectories influenced by transition probabilities.
@@ VII Conclusion and Discussion
In this paper, we have studied a decentralized MBRL-based control solution for CAVs with a limited communication range. In particular, we have proposed the MA-PETS algorithm with a significant performance improvement in terms of sample efficiency. Specifically, MA-PETS learns the environmental dynamics model from samples communicated between neighboring CAVs via PE-DNNs. Subsequently, MA-PETS efficiently develops TS-based MPC for decision-making. Afterwards, we have derived UCRL2-based group regret bounds, which theoretically manifests that in the worst case, limited-range communications in multiple agents still benefit the learning. We have validated the superiority of MA-PETS over classical MFRL algorithms and demonstrated the contribution of communication to multi-agent MBRL.
There are many interesting directions to be explored in the future. For example, we have designed MA-PETS based on a known reward function. However, effective reward design is still an important research topic in the field of RL. Moreover, the ideology of PETS (e.g., ensemble and bootstrap) has been leveraged into many MBRL algorithms, and its integration with the latest DNN solution can be investigated to further improve the performance. Finally, we have only derived the worst case of group regret bound, and the results for more general cases can be explored.

### v2

@@ VI Experimental Settings and Numerical Results
VI-A Experimental Settings
In this section, we evaluate the performance of MA-PETS within the domain of autonomous vehicle control and demonstrate the superiority of our proposed algorithm over several other state-of-the-art RL methods, including FIRL [33], SVMIX [13], MAPPO [8], MADDPG [10], DQN [59], SAC [60]. Specifically, we run our experiments using the CAV simulation platform SMARTS [28] and select the “Unprotected signal-free Intersection” scenario in the closed single-lane and multi-lane “Figure Eight” loop, which is a typical mixed autonomy traffic scenario as illustrated in Fig. LABEL:fig3, for the evaluation. In these experiments, we deploy a fixed number (i.e., $I$) CAVs respectively, managed by our MA-PETS algorithm, and a random number $I_{\text{hv}}$ of human-driven vehicles (HVs) controlled by the environment within the SMARTS framework. All CAVs will start from a one-way lane with an intersection, and drive circularly by passing through the intersection while avoiding collisions and congestion. Following the MDP defined in Section III, the corresponding MDP for CAVs in “Unprotected Intersection” is defined as below.
@@ VI-A Experimental Settings
  • Action: As each CAV is controlled via the target velocity $v\in\mathbb{R}$ and the special actions indicating whether to make a lane change $c\in\{0,1\}$, decided by itself, the action of vehicle $i$ can be represented as $a_{t}^{(i)}=\{v_{t}^{(i)},c_{t}^{(i)}\}$. Besides, for the single-lane scenario, $c$ nulls.
  • Reward: CAVs aim to maintain a maximum velocity based on no collision. Accordingly, the reward function $\mathcal{R}^{(i)}$ can be defined as
@@
In our configuration, the number of episodes is $K=15$, and in the event of no collisions, the length of each episode can reach up to $H=200$ for single-lane scenarios and $H=400$ for multi-lane scenarios, respectively. All of our experiments are conducted on the NVIDIA GeForce RTX 4090 with $5$ independent simulations. Moreover, typical parameters, including the simulation environment setup, deep neural network, and MA-PETS, are summarized in Table II.
TABLE II: List of Key Parameter Settings for the Simulation.
Parameters Description Value
Simulator
Discrete-time step $\tau=0.1\mathrm{~{}s}$
Vehicle numbers $I=8$ or $22$
Human-driving Vehicles $I_{\text{hv}}\in[8,20]$
Distance of “Figure Eight” scenario $480\mathrm{~{}m}$
Maximum velocity of “Figure Eight” scenario $v_{\max}=13.89\mathrm{~{}m}/\mathrm{s}$
Minimum velocity of “Figure Eight” scenario $v_{\min}=0\mathrm{~{}m}/\mathrm{s}$
MA-PETS
Speed per vehicle at time-step $t$ $v_{t}\in[0,13.89]\mathrm{~{}m}/\mathrm{s}$
@@
Target velocity per vehicle at $t$ $\bar{v}_{t}\in[0,13.89]\mathrm{~{}m}$
Speed of vehicles ahead and behind $v_{t,a},v_{t,e}\in[0,13.89]\mathrm{~{}m}/\mathrm{s}$
Distance of vehicles ahead and behind $l_{t,a},l_{t,e}\in[0,75]\mathrm{~{}m}$
@@
Maximal length per episode $H=200$ or $400$
Communication range $d\in[0,200]\mathrm{~{}m}$
@@
Accuracy of CEM Optimizer $\epsilon=0.001$
Seed number $5$
Probabilistic Ensemble Neural Network
No. of Ensembles for dynamics Model $B=5$
Hidden layer number $3$
Hidden units number $300$ or $400$
Buffer length $2,048$
Batch size $128$
Epochs $5$
Learning rate $lr=0.001$
Optimizer Adam
  • $\ast$ The parameters setting used only by MA-PETS.
VI-B Numerical Results
@@
Firstly, we evaluate the performance of MA-PETS for CAVs with $I=8$ and communication range $d=100$ in the single-lane scenario, and present the performance comparison with MFRL algorithms in Fig. LABEL:utility_single_lane. Besides, we testify the real-time performance for each step of the $10$-th episode in the simulation process in Fig. LABEL:real-time. It can be observed from Fig. LABEL:utility_single_lane and Fig. LABEL:real-time that our algorithm MA-PETS significantly yields superior outcomes than the others in terms of both agility and safety. More importantly, MA-PETS converges at a faster pace.
Building upon these single-lane results, we further increase the complexity of the simulation scenarios by establishing a multi-lane intersection without traffic lights and utilizing $I=22$ CAVs controlled by our algorithm MA-PETS, as depicted in Fig. LABEL:utility_multi_lane. In this setting, we perform a comparative evaluation of our algorithm against a suite of state-of-the-art RL methods with a communication range $d=100$. As can be observed from Fig. LABEL:utility_multi_lane, while our algorithm may initially exhibit a marginally reduced efficacy in terms of collision avoidance relative to MAPPO, SVMIX, and SAC, it nearly achieves a collision-free state by the $13$-th episode. Simultaneously, our algorithm MA-PETS notably excels in agility and achieves convergence at a significantly accelerated pace compared to other algorithms.
To validate the rationality of Assumption 1 and the feasibility of MA-PETS after quantization, we conduct supplementary simulation experiments. We compare the performance of MA-PETS using the tile coding method [53] to discretize our six-dimensional continuous state space and two-dimensional continuous action space with MA-PETS operating in continuous state and action spaces within a single-lane “Figure Eight” loop scenario, depicted in Fig.LABEL:simulation_discretization. It can be observed that although the discretized MA-PETS algorithm after quantization exhibits larger fluctuations during the learning process, with the increase in the number of training episodes (e.g., post-$15$ episodes), the MA-PETS with discretization can also achieve performance similar to the continuous MA-PETS in terms of agility and safety. This demonstrates that MA-PETS remains effective even in a discretized setting after quantization.
@@
Furthermore, we show the performance of MA-PETS concerning different values of communication range $d$ by varying from 00 to $200$ in Fig. LABEL:fig_d and Table III. It can be observed from Table III that consistent with our intuition, the increase in communication range leads to a significant boost of the communications overheads. Meanwhile, as depicted in Fig. LABEL:fig_d, it also benefits the learning efficiency of the CAVs, thus greatly upgrading agility and safety. On the other hand, Fig. LABEL:fig_d also unveils that when the communication range increases to a certain extent, a further increase of $d$ contributes trivially to the agility and safety of learned policies. In contrast, Table III indicates an exponential increase in the average communications overhead. In other words, it implies a certain trade-off between the learning performance and communication overheads. Consistent with the group regret bound detailed in Section V, we further investigate the numerical interplay between the minimum number of clique covers $\bar{\chi}\left(\mathcal{G}_{d,k}\right)$ and communication distance $d$ in $15$ independent trials. It can be observed from Fig. 7 that while the minimum number of clique covers remains constant beyond a specific communication range threshold, a generally inverse correlation exists between the minimum clique cover number and the communication radius, which aligns with both our theoretical proofs and the experimental findings presented in Fig. LABEL:fig_d.
@@
Communication Distance $d$ 00 $50$ $100$ $150$ $200$
Communication Overheads 00 $330.13$ $1,866$ $6,139.7$ $10,961.06$
@@
To further illustrate the feasibility and robustness of MA-PETS under more complex and variable real-world communication scenarios, such as vehicles experiencing communication hindrances while passing through tunnels or other challenging environments, we conduct simulations under dynamic and unpredictable communication constraints. In our simulations, we assume the possibilities of each CAV controlled by the MA-PETS to potentially encounter communication blockages or failures as $0\%$, $25\%$, and $50\%$ when attempting to communicate with other vehicles within a communication range of $d=100$ at the end of each episode. The results of these simulations are presented in Fig. LABEL:com_constraint. From Fig. LABEL:com_constraint, it is evident that dynamic and unpredictable communication constraints at varying probabilities can cause fluctuations in the MA-PETS’s safety and agility to some extent. However, the algorithm can effectively learn within $15$ episodes, demonstrating its feasibility and robustness in real-world autonomous vehicle control environments.
To clarify the effect of the MPC horizon $w$ in MA-PETS, we perform supplementary experiments. As depicted in Fig. LABEL:fig_w, from a security point of view, extending the MPC horizon $w$ gradually enhances security. However, beyond a horizon length of $w=25$, further increases yield only marginal security improvements. In terms of agility, the choice of the MPC horizon $w$ is crucial for the performance of the system. In particular, a “too short” horizon introduces a substantial bias in the performance of MA-PETS, as it hampers the ability to make accurate long-term predictions due to the scarcity of time steps. Conversely, a “too long” horizon also results in a noticeable bias. This is due to the increased divergence of particles over extended periods, which reduces the correlation between the currently chosen action and the long-term expected reward. This bias escalates as the horizon $w$ increases. Hence choosing an excessively long horizon adversely affects performance.
Our research also investigates the influence of the number of ensembles $B$ and the number of particles $P$ in MA-PETS. In terms of agility, Fig. 9 and Fig. 10 show the corresponding results respectively, which are derived from $15$ training episodes through $15$ independent simulation runs. It can be observed from Fig. 9 that along with the increase in $B$, the learning becomes more regularized and the performance improves. However, the performance improvement is no longer apparent for a sufficiently large $B$, and this improvement is more pronounced in more challenging and complex environments that require the learning of intricate dynamical models, leaving more scope for effective exploitation of the strategy without the use of model integration. In Fig. 10, superior performance can be reaped for a larger number of particles, because more particles allow for a more accurate estimation of the reward for state-action trajectories influenced by transition probabilities.
@@
To conduct a more in-depth numerical analysis of the complexity and computational time of MA-PETS, we compare the average decision-making time and its standard deviation of different MARL methods after learning 15 episodes, in both single-lane loop and multi-lane loop scenarios as shown in Table IV. The results reveal that the average decision time of MA-PETS is slightly higher than that of FIRL and MAPPO in both scenarios, but faster than SVMIX. Overall, compared to FIRL and MAPPO, MA-PETS has smaller variance and lower variability in decision times, demonstrating greater stability.
[…]
Scenarios The Different MARL Methods
MA-PETS FIRL MAPPO SVMIX
Single-lane loop $8.851\pm 0.3105$ $7.604$\pm$0.3034$$ $8.007$\pm$0.5486$$ $10.28$\pm$0.2668$$
Multi-lane loop $9.726$\pm$0.2615$$ $8.104$\pm$0.4519$$ $8.779$\pm$0.4607$$ $11.06$\pm$0.3025$$
@@ VII Conclusion and Discussion
In this paper, we have studied a decentralized MBRL-based control solution for CAVs with a limited communication range. In particular, we have proposed the MA-PETS algorithm with a significant performance improvement in terms of sample efficiency. Specifically, MA-PETS learns the environmental dynamics model from samples communicated between neighboring CAVs via PE-DNNs. Subsequently, MA-PETS efficiently develops TS-based MPC for decision-making. Afterward, we have derived UCRL2-based group regret bounds, which theoretically manifests that in the worst case, limited-range communications in multiple agents still benefit the learning. We have validated the superiority of MA-PETS over classical MFRL algorithms and demonstrated the contribution of communication to multi-agent MBRL.
There are many interesting directions to be explored in the future. For example, MA-PETS encounters significant challenges in more realistic scenarios. As environmental complexity escalates, scenarios with numerous variables or agents heighten the learning and computational requirements of MA-PETS. Moreover, if the dynamics of the environment are unpredictable, MA-PETS may face difficulties due to the probabilistic nature of its ensemble predictions. Another critical factor is the availability of computational resources. The computational power at hand can significantly influence the performance of MA-PETS, especially as more intricate environments or accurate ensemble methods demand robust processing capabilities. The key next steps involve optimizing the algorithm, reducing its reliance on computational resources, minimizing its decision delays, and improving its adaptability to rapidly changing and complex environments. Furthermore, MA-PETS confronts significant OOD challenges resulting from scarce training data, which can lead to learning instability and substantial overhead. Consequently, designing more accurate models to mitigate OOD issues is a critical priority that requires urgent attention. Finally, we have only derived the worst case of group regret bound, and the results for more general cases can be explored.